If X=(-1,2,0), Y= (3,1,2), Z= (4,-1,0),show that linear combination at (0,1,-1).


1
Expert's answer
2021-08-19T17:23:04-0400
"W=aX+bY+cZ"

"\begin{matrix}
 -a+3b+4c=0 \\
\\
 2a+b-c=1 \\
\\
 2b=-1 \\
\end{matrix}"

"A=\begin{pmatrix}
 -1 & 3 & 4 & & 0 \\
 2 & 1 & -1 & & 1 \\
 0 & 2 & 0& & -1 \\
\end{pmatrix}"

"R_1=-R_1"


"\begin{pmatrix}
 1 & -3 & -4 & & 0 \\
 2 & 1 & -1 & & 1 \\
 0 & 2 & 0& & -1 \\
\end{pmatrix}"

"R_2=R_2-2R_1"


"\begin{pmatrix}
 1 & -3 & -4 & & 0 \\
 0 & 7 &7 & & 1 \\
 0 & 2 & 0& & -1 \\
\end{pmatrix}"

"R_2=R_2\/7"


"\begin{pmatrix}
 1 & -3 & -4 & & 0 \\
 0 & 1 &1 & & 1/7 \\
 0 & 2 & 0& & -1 \\
\end{pmatrix}"

"R_1=R_1+3R_2"


"\begin{pmatrix}
 1 & 0 & -1 & & 3/7 \\
 0 & 1 &1 & & 1/7 \\
 0 & 2 & 0& & -1 \\
\end{pmatrix}"

"R_3=R_3-2R_2"


"\begin{pmatrix}
 1 & 0 & -1 & & 3/7 \\
 0 & 1 &1 & & 1/7 \\
 0 & 0 & -2 & & -9/7 \\
\end{pmatrix}"

"R_3=-R_3\/2"


"\begin{pmatrix}
 1 & 0 & -1 & & 3/7 \\
 0 & 1 & 1 & & 1/7 \\
 0 & 0 & 1 & & 9/14 \\
\end{pmatrix}"

"R_1=R_1+R_3"


"\begin{pmatrix}
 1 & 0 & 0 & & 15/14 \\
 0 & 1 & 1 & & 1/7 \\
 0 & 0 & 1 & & 9/14 \\
\end{pmatrix}"

"R_2=R_2-R_3"


"\begin{pmatrix}
 1 & 0 & 0 & & 15/14 \\
 0 & 1 & 0 & & -1/2 \\
 0 & 0 & 1 & & 9/14 \\
\end{pmatrix}"

Then "a=\dfrac{15}{14}, b=-\dfrac{1}{2}, c=\dfrac{9}{14}."



"\begin{pmatrix}
 0 \\
 1 \\
 -1 \\
\end{pmatrix}=\dfrac{5}{14}\begin{pmatrix}
 -1 \\
 2 \\
 0 \\
\end{pmatrix}-\dfrac{1}{2}\begin{pmatrix}
 3 \\
 1 \\
 2\\
\end{pmatrix}+\dfrac{9}{14}\begin{pmatrix}
 4 \\
 -1 \\
 0 \\
\end{pmatrix}"